Grade 9General Physics

The velocity of the wind over the surface of the wing of an aeroplane is 80mps and under the wing 60mps. If the area of the wing is 8m2 , the dynamic lift experience by the wing ( density of air is 1.3kg.m-3)

Sneha
7 Years agoGrade 9

1 Answer

Vikas TU
7 Years ago
From Bernoulli’s equation,
P1 + 0.5pv1^2 = P2 + 0.5pv2^2
P1 – P2 = 0.5p(v2^2 – v1^2) => 0.5*1.3*(80^2 – 60^2) => 1820
 
Lift = Preseure*A => (P1 – P2)*A => 1820*8 => 14560 N.